A circular loop in the plane of a paper lies in a 0.45 T magnetic field pointing into the paper. The loop's diameter changes from 21.0 cm to 6.2 cmin 0.54 s .

Part A

What is the direction of the induced current?

(clockwise or counterclockwise)

Part B

What is the magnitude of the average induced emf?

Express your answer to two significant figures and include the appropriate units.

| E_avg| = ?

Part C

What is the average induced current if the coil resistance is 4.1 Ω ?

Express your answer to two significant figures and include the appropriate units.

I= ?
Answer #1

here,

magnetic feild , B = 0.45 T

initial diameter , d1 = 21 cm = 0.21 m

final diameter , d2 = 6.2 cm = 0.062 m

time taken , t = 0.54 s

a)

as the magnetic feild is pointing in to the paper

and the flux is decreased,

the induced current is so that the magnetic flux through the loop is increased

so, the current induced is clockwise

b)

the magnitude of average induced emf , EMF = B* pi * ((d1/2)^2 - (d2/2)^2 ) /t

EMF = 0.45 * pi * ( (0.21/2)^2 - (0.062/2)^2) /0.54

EMF = 0.026 V

c)

the average induced current if the coil resistance , I = EMF/t

I = 0.026 /4.1 = 6.34 * 10^-3 A